Latest News/Problems

Jump to: navigation, search

Report here up-to-date issues or problems (and report on any updates to the issue). Do also check the Observer's Log.

March 2014

STX camera sent to manufacturer, to fix faulty shutter. STL 1001E camera mounted as temporary replacement.

September 2013

  • 23-09-2013
Router replaced, connection to control PCs restored, but PIRATE ePowerSwitches still offline 
  • 09-09-2013
PIRATE and PTST offline after lightning strike on 09/09/2013 destroyed main OAM router. 

October 2012

Full control system upgrade with new control PC (July 2012) completed. 
Running Windows 7, ACP 6 and TheSkyX (TSX), all auxiliary software updated.
Facilit is fully operational, but automated pier flip from within ACP remains unreliable
(a known issue in TSX, currently being addresed by SOftware Bisque and DC3).

June 2011

New main imaging camera fitted (SBIG STX-16803).
New control PC installed; stability problems are being diagnosed.
PTST dome took delivery of OTA and mount. Software installation continues.

April 2011

  • 03-04-2011
New PIRATE dome UPS and segment motors installed.
  • 1-04-2011
PIRATE partner observatory PTST 4.5 m dome assembled on OAM west tower. Awaiting 
a telescope!

March 2011

  • 15-03-2010
After ongoing still undiagnosed PC problems finally decided to decommission 
the control PC and re-connect the previous one.

January 2011

  • 27-01-2011
New control PC connected. Awaiting software installation.

August 2010

  • 18-08-2010
PlaneWave CDK17 (PIRATE Mark 2) operational. The Celestron-14 OTA is 
in safe storage for now.

February 2010

  • 20-02-2010
The problems we were having with network switches 
(not set up the network correctly after power failure)
should be a thing of the past when Xisco get some 
replacement switches with a newer firmware this should 
happen some time next week .  
  • 18-02-2010
Yesterday evening Juan and Antonio did fine adjustments
to the Motor 4 - Dome cogwheel interface so that it runs 
smoothly. Dome opens and closes successfully. 
I tried it once this morning. 

The dome should not be opened and closed excessively 
until we get the new motors and dome electronics at latest 
mid march.

But until we then can work kind of normally 
but keep in mind:

leave enough time (2-3 hrs) between closing the dome and the next 
rain front heading towards the observatory so that in case of emergency 
we could alarm someone to close the dome  manually in time. 
Juan and Jaume also found that the Dome UPS had ceased work.
Upon closer inspection Jaume found that a triac had blown (over heated) 
and that all battery cells were emptied below the recharge level.
OAM is sending it back to Germany for repairs under guarantee.
  • 17-02-2010
Dome opens and closes successfully. 

Yesterday Jaume added a spring to Motor 4 to prevent it
from moving as freely as it was ... so that is takes the same
effort to rotate the motor axis by has as for motor 1

We are still trouble shooting but we can work :-)

  • 05-02-2010
One segment motor failed and needs to be replaced.
Baader Planetarium have no motor in stock - PIRATE 
is likely to be out of action for 2-3 weeks.

December 2009

  • 23-12-2009
Juan and Jaume did maintanence today:

 -- weather station: wind and rain measurements working again!
 -- Main CCD camera was taken off the scope, 
            + all filters were cleaned
            + entrance window cleaned 
            + after closing it again all edges were 
              sealed to stop light coming in 
-- The camera it is back on Scepe and connected ...
   fine alignment of the camera was not possible 
   as the dome had to remain closed due to high wind.
  • 18-10-2009

All-sky camera disconnected; awaits repair of USB connection. ENE webcam disconnected, awaits safe enclosure.

November 2009

  • 12-11-2009

All systems running. Both PCs working normally. New dome driver installed.

October 2009

  • 13-10-2009

PC-2 not accessible. Suspect major hardware fault. May need replacement.

  • 12-10-2009

16:00 BST WIMAX working again ...

On 09/10 WIMAX was back swiftly, then off again in the evening due to maintenance (if only this could be announced here!), then ok throughout the weekened. Just now (15:15 BST) it is off again. We really do need a backup link.

  • 09-10-2009

15:35 BST WIMAX connection lost; currently no internet connection to PIRATE.

  • 07-10-2009

20:00 MEST after substituting the RS232 IC we again can connect to the DOME

Now we are checking the PCI to serial boards ... changing the PCI to serial board from PC-2 to PC-1

  • 06-10-2009

15:00 MEST = 14:00 BST = 13:00 GMT

Martin suspects that the RS232 IC i broken ... Jaume has taken the electronic board out and will substitute it today ...

You should be able to go on testing things ... only thing is no remote no manual dome for now ... as the Jaume had to take the littel box with the dome processor out to do the repair.

  • 04-10-2009

we seem to have a communication problem between PC1 and the Dome. they don't seem to establish a connection ... Rebooting PC1 doesn't help.

Also after everything goes off (no more fuel) and is switched on again: no luck ...

On Sept. 26 we left it all running fine ... But bad weather and lack of internet did not let us try it again until last night.

  • 02-10-2009

Observatory is back online with high speed WIMAX link!!

VNC is working.

Tomorrow all other needed ports will be configured

September 2009

  • 30-09-2009

Today at 16:20 hours, has once again lightning has stuck in the OAM, and the router of the phone company has been damage. Tomorrow the phone company technician will como to repair it.

  • 29-09-2009

Guider camera is focussed. Weather sensor reconnected to PC-2. New weather data interface available when power is on (for observers only).

  • 18-09-2009

After earlier lightning strike the OAM automated generator control is disabled. The generator will in general run throughout the night, but will be off during the day to conserve fuel. PIRATE is inaccessible when the generator is off.

The outside webcam awaits a new router.

The guiderscope is not in focus after the camera was removed for use in SXR208. Awaiting clear spell and local help to foucs camera.

  • 01-09-2009
All connections are working again http, ftp, vnc, and the internal webcam ... 
only the external webcam and the allsky image are not running yet.

August 2009

  • 31-08-2009

PIRATE should be reachable again tomorrow as telefonica should have the DSL connections and Xisco hopefully gotten around to reconfigure the ports on the routers

  • 26-08-2009

PIRATE is offline due to lightning strike at 15:00 MT DSL routers all destroyed ... PIRATE computers are all ok

  • 13-08-2009

PIRATE is now in its new dome. While installation work continues most links on PIRATE webpage do not work.

  • 02-08-2009

PIRATE to move into its new dome in the week 10/8-16/8.

June 2009

  • 25-06-2009

PIRATE observer team on site until 3 July 2009, for hardware tests.

April 2009

  • 18-04-2009

Weather sensors operational again.

  • 08-04-2009

All weather systems are down due to maintenance. Due back online "within a few days". PIRATE PC was rebooted to allow joystick operation.

March 2009

  • 05-03-2009

PIRATE wiki software will be upgraded on 6 March, 10:30-11:00 a.m.

February 2009

  • 17-01-2009

Cloud sensor graph working again.

PIRATE wiki will be unavailable for 1 night this week for a software upgrade. We do not know which night.

Cloud sensor graph is stuck at 13/02/2009. Requested restart. While theis is unavailable please use live cloud sensor data available via the link on the PIRATE Observing page.

January 2009

  • 27-01-2009

Burwitz 19:01, 27 January 2009 (GMT) Allsky camera back online weather data working...

Weather data back online. All systems working.

PIRATE PC was rebooted. OAM Roof-top webcam operational again. Still no connection to weather data. Proceed with observations only if satellite image is unambiguous.

  • 26-01-2009

All communications with weather and sky condition sesnors lost. Reboot of router requested.

  • 14-01-2009

Control PC required reboot via remote switch. Camera back on.

  • 13-01-2009

Camera removed for baking of disseccant.

  • 07-01-2009

All systems working, but increased risk of dew formation on chip. Requested baking of dissecant cartridge. Original ACP scripts have been re-installed to avoid frame clutter in MaximDL.

December 2008

  • 17-12-2008

No response from devices with cloud sensor IP address - again.

Rooftop webcam back on-line. ACP restarted.

  • 15-12-2008

Weather sensor, cloud sensor and all-sky camera online again. Rooftop webcam still off.

  • 12-12-2008

Weather sensor, cloud sensor, rooftop webcam and all-sky camera are currently off-line.

I've changed two of the ACP scripts to use fits extension instead of fts. Also changed Maxim to use fits. Copies of the unchanged scripts are in folders called Backup in the ACP root folder and ACP\Scripts folder. Rob.

  • 11-12-2008

PC was rebooted, panyc restarted and functions verified, ACP restarted, connected and disconnected. PIRATE should be ready for use.

November 2008

  • 28-11-2008

PC appears to have rebooted itself today. All-sky camera is apparently stuck. Camera cartridge was allowed to dry out in main observatory building; cartridge presumably baked again.

  • 8-11-2008

PIRATE dehumidifyer is now running 5-10 hrs/day, depending on generator use. Camera cartridge was baked for 4 hrs at 180° C and replaced in camera. Schmidt plate was cleaned.

  • 6-11-2008

PIRATE enclosure protective seal and dehumidifier installed. Dew system fully operational. Camera cartridge to be baked.

  • 3-11-2008

All weather systems fully functioning now, yet to hear back from the OAM regarding the functioning (?) of the dew system (fans & heater strips), so have chased up on this.

October 2008

  • 30-10-2008

Weather systems down as of ~ 3:am. All-sky camera down, weather station also stopped updating, Cloud sensor down.

PC needs rebooting? Will investigate tomorrow.

  • 24-10-08

Microsoft security update triggered automatic reboot of control PC. Relay needs restart - wait for rain to clear.

Relay working again.

  • 22-10-08

All systems up and running again. The ADSL failure was a wider telephone line problem in the area.

  • 21-10-08

Complete data link black out - no access to any PIRATE component; suspect OAM us cut off from internet.

  • 19-10-08

Cloud sensor and all sky cam functioning again.

  • 17-10-08

Cloud sensor is also not working, along with the all-sky camera. The weather station appears to still be working.

  • 16-10-08

All-Sky camera is not updating, and is stuck on an image from 14/10/08

  • 15-10-08

PC 2 was rebooted, and the security upgrade applied. As a result, ACP is no longer running. Can PC downtimes please be announced on the wiki in advance!

  • 02-10-08

New bulb was fitted, PC rebooted. All systems go at the moment ...

  • 01-10-08

Asked for a new bulb to be fitted today.

September 2008

  • 30-09-08

PIRATE PC was non-responsive; remote log-in not possible; PC rebooted by OAM

PIRATE lamp needs a new bulb. Currently the light is an essential tool - it is the only independent way to check if PIRATE is in the park position, and if it is safe to close the roof.

Await decision by Rob if Windows Service Pack can be installed.

  • 29-09-08

PIRATE light doesn't work (panyc signals light is on, but webcam shows it isn't) - can't use PIRATE without light at night as roof status can't be verified

Not clear if power is off to other devices as well. Powering up devices is not recommended with roof closed, as there is always a risk that PIRATE attempts to slew to home position accidentally. Check roof by daylight.

This clearly demonstrates the need for redundancy (2nd light), an IR webcam, and a communication channel to OAM.